OpenStudy (anonymous):

-3x+3y=3, 3x+y=9 please solve by using the elimination method

OpenStudy (anonymous):

-3x+3y=3, 3x+y=9 please solve by using the elimination method

OpenStudy (tkhunny):

Add them. -3x + 3x = 0 3y + y = 4y 3 + 9 = 12